OpenStudy (anonymous):

AP Calculus: Slope Fields/ Separate Variables Find the equation of the curve, given the derivative and the indicated point on the curve. dy/dx=1/(x+2)

OpenStudy (anonymous):

point on graph is (0,1)

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (tkhunny):

Can you find the general antiderivative of 1/(x+2)?

OpenStudy (anonymous):

u mean integral?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

You can call it "Steve" if you want. That's what you need to solve it. What do you get?

OpenStudy (anonymous):

soo far this is what I got \[\int\limits_{}^{} dy = \int\limits_{}^{} \frac{ 1 }{ x+2 } dx\] \[y=\ln \left| x+2 \right| + C\]

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Perfect. Why didn't you do that in the first place? Maybe you did and you just weren't sharing. Okay, now substitute the known initial values and solve for "C". x = 0 y = 1

OpenStudy (anonymous):

how do i get rid of ln though

OpenStudy (tkhunny):

Why do you want to do that? Are you a logarithmist? Just solve it with whatever it is.

OpenStudy (anonymous):

ok so i got \[1=\ln \left| 2 \right| + C\]

OpenStudy (tkhunny):

2 > 0 Simplify and solve for C. You'll be done.

OpenStudy (anonymous):

2>0? I still dont get it how do i get C by itself?

OpenStudy (anonymous):

wouldn't i times ln by e to cancel out "ln"? @tkhunny

OpenStudy (anonymous):

or put in a \[ \pm\]

OpenStudy (tkhunny):

That makes absolutely no sense at all. It seems as though you are throwing dice to try to solve this. 1) Never use "times" like that again. It's multiplication. Multiply things. 2) Just solve for C. Why are you resisting? 1 = ln(2) + C C = 1 - ln(2) Why isn't that done? Is there a federal mandate to banish all logarithms?

OpenStudy (anonymous):

ok ty :)

OpenStudy (anonymous):

@tkhunny can you help me with another one?

OpenStudy (tkhunny):

Only if you post it and promise to quit hating on logarithms. ;-)